let y > 0 be a regionof spacewith a uniform and constant magnetic feild Bk^ . a particle with charge q and mass m travels along the y axis and enters into a magnetic feildat origin with speed v0 . in the region the particle is subjested to an additional friction force F=-kv . assume that particle remains in the region y > 0 at all times . the coordinates of the particle where it will finally stop are (x,y) then x=?  ,y=?  

Grade:12th pass

1 Answers

ROSHAN MUJEEB
3 years ago
SIMPLY apply kinematic equation like a=dv/dt, a=vdv/dx, v=dx/dt etc